OpenStudy (goformit100):

Which is greater 3^00 or 2^300

OpenStudy (goformit100):

@noided

OpenStudy (mathstudent55):

What is 00?

OpenStudy (anonymous):

2^300

OpenStudy (anonymous):

3^200 or 2^300 ?

OpenStudy (nathan917):

2300 = (23)100 = 8100 3200 = (32)100 = 9100 and since 9 > 8, 3200 is greater than 2300

OpenStudy (anonymous):

its 3200

OpenStudy (goformit100):

can we do this problem mentally ?

OpenStudy (anonymous):

a^0=1

OpenStudy (anonymous):

purely mathematically the number 00 doesn't really exist, so I guess you meant 0?

OpenStudy (anonymous):

3^200=9^100 2^300=8^100 exponents are the same compare only base 9>8 ==> 3^200> 2^300

OpenStudy (goformit100):

ok

OpenStudy (goformit100):

Thank you sir.

OpenStudy (anonymous):

np